OpenStudy (anonymous):

the number of arrangements can be made taking 4 letters at a time out of the letters of the word PASSPORT

OpenStudy (dan815):

P,A,S,O,R,T P S

OpenStudy (dan815):

case1) no repeating letters case2)repeating Ps only case3)repeats S only case4)repeating P and s

OpenStudy (dan815):

case1)=6! case2)=5*4*(4choose2) case3) =5*4*(4choose2) case4) = (4 choose 2)

OpenStudy (anonymous):

In the case one, should't it be 6 x 5 x 4 x 3 and not 6 ! and can you please explain the last case?

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

that should be 6*5*4*3 yes thats a mistake

OpenStudy (dan815):

the last case we are looking at P,P,S,S how many ways to arrange that

OpenStudy (dan815):

thats 4!/(2!2!)

OpenStudy (anonymous):

And in case 2 and 3 it will be like : 5 x 4 x 4!/2! x 2! ? Why 4!/2! x 2! ? ( that will be similar as case 4 then? @dan815

OpenStudy (dan815):

yeah because

OpenStudy (dan815):

uh its too much to explain typing =.=

OpenStudy (dan815):

just think about it lol